OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty-five dollars. This year, he paid one hundred sixty-two dollars. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

I think 20% can you check my work? 162-135=27 27/135=.2 .2*100=20 20% Soooooo....

OpenStudy (anonymous):

Yep, you're method is spot on! Well done!

OpenStudy (anonymous):

thanks